please with steps


NH4NO2 (s) = N2 (g) + H2O (g)

When a sample is decomposed in a test tube, 949 mL of wet N2(g) is collected over water at 26°C and 778 torr total pressure. How many grams of dry NH4NO2(s) were initially decomposed? The vapor pressure of water at 26°C is 25.2 torr.

Mass=.............g




1
Expert's answer
2021-03-09T05:55:20-0500

1.find the amount of substance

"PV = nRT"

P=778-25,2=752.8=0.990526712 atm

26°C=299.15 K

"0.990526712\times0.949 = n\times0.08206\times299.15"

"0.94 = n\times24.55"

n=0.0383

"m=M\times n=64.044\times0.0383=2.45 g"